Summary

Note that the pointing was off by ~5 arcmin, and we discovered this at 2am while tuning things up for the Yan program. Seeing was unstable the whole night from 1.5" - 2.3". We got Milne data and everything else I will be suspicious. Vilas comet images had better quality seeing from 1.5-1.7". Comet was not seen in a single image but the nominal position was next to the big star....so this may be the issue.

Details

Observers: David Sand and Jana Pittichova

Supporting Astronomer: Marco Pedani

Operator: Wayne Wack

All Times UT

Second Night of the AZ queue.

2350 Sky looks OK. Lots of high cirrus. The mirror and ambient temperatures are almost identical. BOOYAH.

???? Begin taking biases

0120 We stopped biases to let the day crew rebalance the telescope.

0133 Be careful with biases. Every once in a while a red bias will show up saturated. Weird.

0140 Wind is at 15 m/s pretty constant, but rising slowly.

0200 They are having trouble getting both shutter doors open. I worry we will not be able to take sky flats.

0221 Moved to Segue_0_bino. There is a y-limit on one of the mirrors. Wayne will fix it, and we'll focus.

0226 focusing.

0253 still focusing.....my god

0303 Red seeing is 1.5 and blue is 1.8, according to focus. We will try to do Segue anyways.

0304 Segue_0_bino -- seeing is 2 arcseconds in blue and red. we're going to move to Milne's bad seeing field. Note that Segue_0_bino is NOT done. Seeing is 2 arcsec.

0318 Moving to SN2008D to do milne_sn2008D_BR

0320 Focus

0336 milne_sn2008D_BR -- Seeing is about 1.3 arcsec in blue and 1.5 in red. Neither has very round images, though. Will return to Segue after this OB. Wind still about 10 m/s Seeing is 1.2 arcsec in second image. Seeing is improving. Seeing is 1.6 arcsec in third image. Images are a bit elongated, because we are nearing zenith. We will finish this OB though Last image has some tracking issues since we are pointed very close to zenith.

0401 Moving back to Segue_0_bino. I hope seeing is better here.

0412 restarting Segue_0_bino. John Hill, Olga and I notice a big guide star jump on the red side. Seeing is about 1.7 arcsec on both sides for first two image. BLURG. I'll take it. Seeing is getting worse. I will finish this OB, but then I'm going back to the bad seeing program. 0438 Seeing is now 2 arcseconds. I'm stopping the OB. Sorry to waste time, this is just no longer worth it.

0443 Moving to do milne_sn2008D_VI. Focusing.

0453 beginning milne_sn2008D_VI -- First image seeing is 1.3 arcsec in blue and 1.2 in red. I hate going back and forth. Second image -- 1.4 in blue; 1.3 in red Third image -- 1.6 in blue; 1.55 in red Fourth image -- 1.5 in blue; 1.55 in red and images look funny

0520 I bet Y band images will be fine, so we are moving to b63r.

0542 We took 2 images for b63r, seeing is around 2" for both red and blue, going back to the bad seeing program

0551 Focusing and than Milne SN2007sr program

0612 first image of SN2007sr_BR, seeing is 2.3" for the blue and 1.93" for the red, seeing getting worst for red channel

0625 temperature changing and warming up to soon, 3C in 30 minutes

0700 we lost 30 minute to focus...there was problem with blue channel because 2 big stars were in the field and the dofpia was not able to fit the background. No problem for red channel. Finally we had to offset the telescope.

0712 after long time trying to focus blue channel we get somehow good focus and will start SN2007sr_VI Both blue and red images look pretty bad. as much as it pains me, I will recollimate. There may be a little bit of cirrus. Number counts on the polaris monitor are down.

0730 retrying milne_sn2007sr_VI Both images a bit elongated. John Hill and I can't figure it out-- the collimation seems fine. We will finish the OB. seeing roughly 1.6 arcsec in the red seeing is getting better. you can see it in the polaris monitor. 3rd red image is 1.3 arcsec.

0754 Seeing has improved. Moving to boot63. Go science. Focusing.

0817 beginning boot63 -- First image -- U image has no guide star. Y is guiding on a big fat galaxy. U image has 1.5" seeing. Y is bad probably because it was guiding on galaxy. Fuyan, check the images in this OB before you consider it done. Several of the individual images were guiding on galaxies. When we guided on stars, the image quality was ~1-1.5 arcsec.

0842 Moving to Yan field. OSUnb01

0916 Spent last half hour checking the pointing. Jesus. Pointing was 5 arcmin off ALL NIGHT

0919 beginning focus.

0931 OSUnb01 begun. Narrow band is funny shaped. I doubt you can use it without guiding. Sand has decided we should move to the comet.

0945 Move to the comet 103P/Hartley 3 and focus, start for red channel are elongated and seeing is 1.78". stars on the blue channel are OK and seeing is 1.96"

1100 Possible comet candidate visible on single image....but just next to a big star and is hard to tell if it is really moving object. smile Seeing is little worst 1.71"

1125 I am not sure about the comet - it can be there but is not very clear...have to co-add images

1200 Seeing is still around 1.57" for blue and 1.76" for red, strange same saturation for one of the comet image as we had for one bias

1214 testing the sky twilight flats VR

1234 twilight sky flat BI

1256 5 biases -- XiaohuiFan - 27 Mar 2009
Topic revision: r4 - 28 Mar 2009, XiaohuiFan
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback